# An Survey of the Evolution of Language Model-Based Dialogue Systems

### v1

@@ 4.2 Fusion between TOD and ODD
The ongoing pursuit of task-level unification in the field of dialogue systems remains a prominent focus. There are lots of work which have embarked on an exploration of integrated task scenarios in order to achieve a more faithful emulation of human-level conversations, where a diverse array of information is seamlessly integrated [8]. This endeavour extends beyond conventional single dialog types, aiming to integrate ODD and TOD to elevate the authenticity and engagement of dialogue interactions [134, 167, 168, 169, 170]. This kind of mix-type dialogue relax the strong assumption that the user must have a clear and specific goal in TOD, enabling the user to clarify their goals through dynamic interactions [8]. Table III shows the high-level fusion of TOD and ODD.
To fuse different data and tasks between TODs and ODDs, there are different ways according to various types of ODDs: social chit-chat, and knowledge-grounded chit-chat. First of all, several methods attempt to build a unified data schema first while saving some important sub-tasks in TODs and formulating open-domain dialog as end-to-end response generation task [134, 167, 170, 8]. Specifically, both UniDS [134] and PLATO-MT [8] choose to optimize three sub-tasks: belief state generation, system act generation, and response generation, in an end-to-end manner. The data schema usually contains (user input, belief state, DB result, system act, system response). In addition, some of the works consider the transition from TOD to ODD [170] while others model the opposite direction: ODD to TOD [168], both aiming to intertwine these two dialogue modes together seamlessly in the same conversation. Specifically, Tom et al. [170] further feature inter-mode contextual dependency between TOD and ODD, i.e., the dialogue turns from the two modes depending on each other. However, the ODD in these unified dialogue systems is mainly for chit-chat, such as “you’re welcome” and “sounds good to me”, which does not require any external knowledge sources [169].
To incorporate more complicated open-domain situations, another line of work considers knowledge-grounded chit-chat [88, 169], enabling dialog systems to access external knowledge so that they can provide more engaging and knowledgeable conversations and reduce hallucinations simultaneously [171]. Some of them propose to insert additional turns into TOD, where the system needs to respond based on the knowledge from domain FAQs [172] or unstructured documents [173]. These datasets mentioned primarily address the task of identifying conversational turns that necessitate external knowledge for generating responses and selecting the pertinent knowledge for response construction. In contrast, KETOD aims to enhance the authenticity and engagement of dialogues by integrating knowledge-grounded casual conversations into the existing TOD responses [169].
@@
TABLE III: High-level of fusion between TOD and ODD, wherein there are different types of ODD.
Model Task Type of ODD Knowledge sources Training method
Database Web PLMs
Q-TOD [135] TOD End-to-end
Internet-Augmented [88] ODD KG Chit-Chat End-to-end
UniDS [134], FusedChat [170] TOD + ODD Chit-Chat End-to-end
SalesBot [168]
HyKnow [173], KETOD [169] TOD + ODD KG Chit-Chat End-to-end
GODEL [153] TOD + ODD Question Answering, KG Chit-Chat End-to-end
DuClarifyDial [8] TOD + ODD Chit-Chat, Question Answering, KG Chit-Chat End-to-end
MDS [174] TOD + ODD Question Answering End-to-end
OPERA [160] TOD + ODD Question Answering End-to-end
DialogStudio [133] TOD + ODD Dial-Sum, NLU, Conv-Rec, KG Chit-Chat End-to-end
@@
Furthermore, more and more works fuse other tasks or skills into the dialogue system. 1) Skills. Madotto et al. [120] introduce the Adapter-Bot, a dialogue model that uses a fixed backbone conversational model such as DialGPT [13] and triggers on-demand dialogue skills (e.g., emphatic response, weather information, movie recommendation) via different adapters, while Smith et al. [175] also try to blend multiple skills into one conversation session, in which knowledge, emotional and personalizing skills are shown together in the same multi-turn conversation. 2) Tasks. Besides the combination of question-answering and TOD tasks mentioned in § 4.1, DuClarifyDial [8] additionally incorporate chit-chat, knowledge-grounded chit-chat while GODEL [153] specifically target knowledge-grounded goal-directed dialogs by leveraging a new phase of grounded pre-training on open-domain goal-directed corpus. There is the latest work, DialogStudio, which further encompasses data from natural language understanding, conversation recommendation, and dialogue summarization [133]. 3) Knowledge Sources. Most of the previous works either rely on the dataset knowledge for TOD or the Web for ODD [135, 88], and recent works generalize the possible knowledge sources to include databases, the Web, and pre-trained language models at the same time [160]. More analysis can be found in § 5.2.
4.3 Fusion between LLM and DM
TABLE IV: An overview of Dialogue Models based on Large Language Models listed in time order.
Dialogue Models Release Date Base Model # of Params Window size Trainig Tokens Language Preference Source Type
BlenderBot3[89] Aug, 2022 OPT 3B/30B/175B 180B EN Open
ChatGPT Dec, 2022 GPT3.5 [176] 175B 4k/16k EN Closed
Bing Chat Feb, 2023 GPT4 [177] 8k/32k EN Closed
LLaMA-Chat [35] Feb, 2023 LLaMA 7B/13B/33B/65B 2k 1.4T EN Open
Alpaca [178] Mar, 2023 LLaMA 7B 2k 1.4T EN Open
Vicuna [179] Mar, 2023 LLaMA 7B/13B 2k 1.4T EN Open
ChatGLM [180] Mar, 2023 GLM 6B/12B 2k 1T CN Open
Slack Mar, 2023 Claude 200k 2.0T EN Closed
Bard Mar, 2023 PALM 540B EN Closed
Falcon-Chat [181] Jun, 2023 Falcon 7.5B/40B/180B 2k 1.5T/1T/3.5T EN Open
MPT-Chat [182] Jun, 2023 MPT 7B/30B 8k 1T EN Closed
InternLM-Chat [183] Jun, 2023 InternLM 7B/20B 8k 2.3T CN Open
Baichuan-Chat [184], Jun, 2023 Baichuan 7B/13B/53B 4k 1.2T CN Open (7B/13B)
ChatGLM2 [180] Jun, 2023 GLM 6B/12B 32k 1T CN Open
LLaMA2-Chat [185] Jul, 2023 LLaMA2 7B/13B/34B/70B 4k 2.0T EN Open
Atom-Chat Aug, 2023 LLaMA2 7B/13B 4k 1.0T CN Open
Qwen-Chat [186] Aug, 2023 QWen 7B/14B 8k 2.4T CN Open
Baichuan2-Chat [184] Sep, 2023 Baichuan2 7B/13B 4k 2.6T CN Open
ChatGLM3 Oct, 2023 GLM 6B 32k / CN Open
@@ 4.3 Fusion between LLM and DM
With the emergence of large language models (LLMs), such as ChatGPT and LLaMA [35, 185], the boundaries between task-oriented and chit-chat dialogues have become increasingly indistinct, since LLMs are beginning to demonstrate a preliminary capability for both task-oriented [187] and chit-chat dialogues [188] even under the zero-shot or few-shot settings. Following the release of ChatGPT, numerous large language models are proposed and corresponding dialogue models are finetuned on different-sized dialogue corpus, aiming to achieve better performance on different dialogue tasks, domains or even languages [34, 185]. Hence, an exponential growth in both LLMs and DMs has been observed. To facilitate a more comprehensive illustration and comparison of the advancements, we here choose to focus on the two most widely used languages, namely Chinese and English as shown in Table IV.
@@
LLMs and DMs in English: There are two major series of LLMs: GPT [11, 25, 177] and LLaMA [35, 185]. On one hand, the evolution of the GPT-series models initiated with GPT [11] and has been consistently refined by OpenAI. This improvement involves scaling both the model size and the pre-training data size, including the incorporation of code data [189] and increasing the model parameters to a larger size, notably reaching 175 billion in GPT-3 [176]. The culmination of these enhancements is evident in the success of ChatGPT, achieved through additional fine-tuning and alignment techniques based on the GPT-3.5 model. Initially trained on a dataset of 570GB, ChatGPT supports input lengths of 4k and 16k, and demonstrates remarkable capabilities in human communication, boasting a substantial knowledge base, proficiency in reasoning through mathematical problems, accurate contextual understanding in multi-turn dialogues, alignment with human values to ensure safe use and following diverse user instructions. Subsequently, the introduction of GPT4 [177] further extends the model’s capabilities, enabling the resolution of multi-modal tasks involving both image and text inputs and outputs. It is worthy noting that the maximum input window size is also doubled, which empowers the model to handle long context input. Benefiting from the cooperation with OpenAI, Microsoft release New Bing  [^6] https://blogs.microsoft.com/wp-content/uploads/prod/sites/5/2023/02/The-new-Bing-Our-approach-to-Responsible-AI.pdf, which is a conversational agent based on GPT4, capabling to provide up-to-date information with the assistance of search engine.
@@
On the other hand, unlike the closed-source nature of the GPT series models from OpenAI, Meta has embraced openness by releasing the LLaMA [35] and LLaMA2 [185], which are pre-trained on 1.4T data, offers four parameter configurations: 7B, 13B, 33B, and 65B. This provides a fresh opportunity for an open-source community, with corresponding dialogue models being consistently finetuned based on LLaMA. LLaMA-Chat, for example, finetunes the LLaMA model using instructional methods, aligning with the GPT-3 autoregressive paradigm. In addition, Alpaca [178] is finetuned by Stanford based on LLaMA-7B, using text-davinci-003 API and self-instruct techniques to create a prompt-response instruction dataset of 52K. Vicuna [179], based on LLaMA-13B, used supervised data from ShareGPT  [^7] https://sharegpt.com/, extending sequence length from 512 to 2048 during training. To further improve the window size and human values alignment with the public, LLaMA2 [185] employs grouped-query attention, extending input length from 2k to 4k and training data from 1.4T tokens to 2.0T tokens. Besides these two series, there are other series English LLMs such as Claude [^8] https://claude.ai/, Bard, Falcon, etc. These models typically adhere to an architecture resembling that of GPT-3, showcasing proficiency in handling multi-turn and complex dialogues, with different sizes of input window. Other commercial companies have also adopted similar training methods and subsequently introduced their own conversational large language models, such as MPT-Chat [182] and Falcon-Chat [181].
@@
LLMs and DMs in Chinese: ChatGLM [34, 36], as the first open-source Chinese LLM, which is a bilingual model (Chinese and English) built on the general language model architecture – GLM 130B [36]. The open-sourced 6B version, ChatGLM2-6B, extends the context length from 2k to 32k tokens and employs multi-query attention for more efficient inference. Furthermore, they progress from ChatGLM and ChatGLM2 to ChatGLM3  [^9] https://github.com/THUDM/ChatGLM3/, which incorporates a broader and more diverse corpus of data during its fine-tuning process. Additionally, ChatGLM3 has demonstrated significant improvements in planning and reasoning, akin to language agents [190, 191], thus reflecting substantial improvements in tool learning, code execution, database operation, knowledge graph search and reasoning. There are publicly available LLMs from other companies, such as Baichuan-Chat [184] and Baichuan2-Chat. The latter, Baichuan2-Chat, employs a method similar to Claude and is trained on over 2.6T tokens. These models cater to a wide spectrum of applications, each offering unique advantages stemming from their fine-tuning strategies and underlying architecture. Alibaba releases Qwen-7B-Chat aligns with human intent using a curated dataset, pre-trained on over 2.2T tokens with a context length of 2048. Sensetime publish InternLM-Chat, which is derived from supervised fine-tuning and reinforcement learning with human feedback on InternLM [183], a multilingual model pre-trained on 2.3T tokens.

### v2

@@ 4.2. Fusion between TOD and ODD
As shown in Table 1, there are different types of dialogue in both TOD and ODD. To seamlessly address them together in an unified manner, lots of work which have embarked on an exploration of integrated task scenarios in order to achieve a more faithful emulation of human-level conversations, where a diverse array of information is seamlessly integrated (Liu et al., 2022b). This endeavour extends beyond conventional single dialog types, aiming to integrate ODD and TOD to elevate the authenticity and engagement of dialogue interactions (Zhao et al., 2022; Sun et al., 2021; Chiu et al., 2022; Chen et al., 2022; Young et al., 2022). This kind of unified dialogue relax the strong assumption that the user must have a clear and specific goal in TOD. Instead, they enable users to refine or discover objectives through multi-turn natural language exchanges, akin to human conversations (Liu et al., 2022b). Table 4 shows the high-level fusion of TOD and ODD. In general, there are two paths to build a unified dialogue system: expanding TOD to incorporate open-ended dialogue or enhancing ODD with task-oriented capabilities.
@@
Table 4. High-level of fusion between TOD and ODD, wherein there are different types of ODD.
Model Task Type of ODD Knowledge sources Training method Database Web PLMs Q-TOD (Tian et al., 2022) TOD End-to-end Internet-Augmented (Komeili et al., 2022) ODD KG Chit-Chat End-to-end UniDS (Zhao et al., 2022), FusedChat (Young et al., 2022), SalesBot (Chiu et al., 2022) TOD + ODD Chit-Chat End-to-end HyKnow (Gao and et al., 2021), KETOD (Chen et al., 2022) TOD + ODD KG Chit-Chat End-to-end GODEL (Peng et al., 2022) TOD + ODD Question Answering, KG Chit-Chat End-to-end DuClarifyDial (Liu et al., 2022b) TOD + ODD Chit-Chat, Question Answering, KG Chit-Chat End-to-end MDS (Nehring and et al., 2021) TOD + ODD Question Answering End-to-end OPERA (Li and et al., 2023) TOD + ODD Question Answering End-to-end DialogStudio (Zhang and et al., 2023a) TOD + ODD Dial-Sum, NLU, Conv-Rec, KG Chit-Chat End-to-end
The first path, extending TOD to ODD, involves taking existing task-oriented dialogue systems and expanding their capabilities to handle more open-ended, free-form conversation. Specifically, since TOD consists of unique data schema such as belief state and database results, some of previous methods attempt to build a unified data schema to making it can extend to open-domain settings, falling into the first category. Lots of other works conceptualize open-domain dialogue as an end-to-end response generation task, where external knowledge is treated as a special type of database result (Zhao et al., 2022; Sun et al., 2021; Young et al., 2022; Liu et al., 2022b). For examples, both UniDS (Zhao et al., 2022) and PLATO-MT (Liu et al., 2022b) choose to optimize three sub-tasks: belief state generation, system act generation, and response generation, in an end-to-end manner. Moreover, Young et al. (2022) further feature inter-mode contextual dependency between TOD and ODD, i.e., the dialogue turns from the two modes depending on each other. To conclude, the unified data schema usually contains (user input, belief state, DB result, system act, system response) following the traditional TOD setting. Instead of proposing an unified schema, a simpler approach is to directly insert additional open-domain dialogues into task-oriented dialogues. In this way, the system responds based on knowledge from domain FAQs (Kim and et al., 2020) or unstructured documents (Gao and et al., 2021) whenever it detects that the current query requires open-domain knowledge. These datasets mentioned primarily address the task of identifying conversational turns that necessitate external knowledge for generating responses and selecting the pertinent knowledge for response construction.
Conversely, the second path, enhancing ODD with TOD capabilities, focuses on equipping open-domain systems with structured and task-oriented interactions. This involves adding the ability to understand user intents, track dialogue state, and complete well-defined tasks, on top of the open-ended conversational capabilities. Previous researchers either use a retrieval-augmented framework (Tian et al., 2022; Peng et al., 2022; Zhang and et al., 2023a), treating external databases (in TOD) or documents (in ODD) as special sources of knowledge, or they modularize each component within the dialogue systems (Li and et al., 2023; Zhang and et al., 2023a). Taking Q-ToD as an example, Tian et al. (2022) introduce a novel query-driven task-oriented dialogue system, namely Q-TOD. The essential information from the dialogue context is extracted into a query, which is further employed to retrieve relevant knowledge records for response generation, following the conventional architecture of open-domain dialogue systems. Similarly, Peng et al. (2022) utilize a similar architecture but focus on the effects of pretraining with a comprehensive dialogue corpus. Furthermore, several attempts apply the modular dialog system framework to combine TOD and ODD. For example, Nehring and et al. (2021) propose a hierarchical architecture composed of several dialogue systems, and then use a classifier to select the right routine to answer current user utterance. We emphasise that both of these two paths share a common objective which is to intertwine these two dialogue modes together seamlessly in the same conversation.
@@
Beyond fusing TOD and ODD, researchers try to incorporate additional skills, tasks, and knowledge sources, enriching the scope and boundaries of dialogue system, making them more adaptable and human-like. i) Skills. Lin et al. (2021) introduce the Adapter-Bot, a dialogue model that uses a fixed backbone conversational model such as DialGPT (Zhang et al., 2020b) and triggers on-demand dialogue skills (e.g., emphatic response, weather information, movie recommendation) via different adapters, while Smith and et al. (2020) also try to blend multiple skills into one conversation session, in which knowledge, emotional and personalizing skills are shown together in the same multi-turn conversation. Some recent studies focus on other skills of dialogue system, such as proactivity (Deng and et al., 2025) and different conversational strategies (Wang and et al., 2023d); ii) Tasks. Besides the combination of question-answering and TOD tasks (as mentioned in § 4.1), DuClarifyDial (Liu et al., 2022b) additionally incorporate chit-chat, knowledge-grounded chit-chat while GODEL (Peng et al., 2022) specifically target knowledge-grounded goal-directed dialogs by leveraging a new phase of grounded pre-training on open-domain goal-directed corpus. There is a latest work, DialogStudio, which further encompasses data from natural language understanding, conversation recommendation, and dialogue summarization (Zhang and et al., 2023a). iii) Knowledge Sources. Most of the previous works either rely on the dataset knowledge for TOD or the Web document for ODD (Tian et al., 2022; Komeili et al., 2022), and recent works generalize the possible knowledge sources to include all databases (i.e., user memory, domain-specific database), the Web, and even pre-trained language model itself at the same time (Li and et al., 2023). More analysis can be found in § 5.2.
The integration of TOD and ODD into unified frameworks offers significant benefits but also introduces new complexities. On the positive side, these systems enhance the user experience by seamlessly interleaving task completion (e.g., booking a hotel) with social interaction (e.g., discussing travel preferences), thereby eliminating the jarring transitions seen in modular systems. This unified approach not only paves the way toward more human-like conversational agents but also improves scalability through joint optimization of diverse sub-tasks. However, the fusion of these dialogue modes presents notable challenges. Unified models require large-scale, complex annotated datasets to accommodate varied data formats and tasks, which significantly increases the cost and difficulty of data curation. Moreover, blending distinct dialogue modalities complicates error analysis, making it more challenging to identify and correct issues. Finally, the inherent trade-off between generality and specificity may lead to performance gaps, where enhancing one modality could inadvertently compromise another, ultimately affecting the system’s overall reliability and interpretability.
4.3. Fusion between LLM and DM
With the continual influx of data, model sizes have grown significantly at the same time (i.e., the scaling law (Kaplan et al., 2020)), leading to the appearance of large language models (LLMs) (Ouyang et al., 2022; Touvron et al., 2023b). These LLMs exhibit exceptional generalization across diverse tasks while minimizing the need for extensive task-specific fine-tuning that traditional dialogue models require. Moreover, the universal training recipe, including instruction-tuning and reinforcement learning from human feedback (RLHF), empower LLMs to adapt flexibly to conversational contexts, bridging the gap between open-domain chit-chat and goal-oriented dialogue.
@@ 4.3. Fusion between LLM and DM
This shift was catalyzed in 2022 by OpenAI’s groundbreaking product — ChatGPT  [^6] https://chatgpt.com/ – a single model that seamlessly integrates contextual understanding, instruction following, and task completion to deliver responses that are both helpful and safe. Consequently, the boundaries between task-oriented and open-domain dialogues have become increasingly indistinct, since LLMs are beginning to demonstrate a preliminary capability for both task-oriented (Zhang and et al., 2023c) and all other types of dialogues (Wang et al., 2023b) even under the zero-shot or few-shot settings. Following the release of ChatGPT, numerous LLMs are proposed and corresponding dialogue models (DM) are finetuned on different-sized dialogue corpus, aiming to achieve better performance on different dialogue tasks, domains or even languages (Du et al., 2022; Touvron et al., 2023b). Hence, there has been an exponential growth in both LLMs and DMs, encompassing both open-source and proprietary (closed-source) models. To provide a more comprehensive analysis and comparison of advancements, we focus on the two most widely used languages — Chinese and English — before discussing multilingual LLMs in the final part. Table 5 shows the overview of dialogue model based on different base models.
@@
LLMs and DMs in English. Out of lots of English models, Meta’s LLaMA series (Touvron et al., 2023a, b) is a leading example of open-sourced LLMs, while OpenAI’s GPT series (Ouyang et al., 2022; OpenAI, [n. d.]) represents one of the most well-known proprietary models. On the one hand, the evolution of the GPT-series models initiated with GPT (Radford et al., 2018) and has been consistently refined by OpenAI. This improvement involves scaling both the model size and the pre-training data size, including the incorporation of code data (Chen et al., 2021) and increasing the model parameters to a larger size, notably reaching 175 billion in GPT-3 (Brown et al., 2020). The culmination of these enhancements is evident in the success of ChatGPT, achieved through additional fine-tuning and alignment techniques from human feedback (Ouyang and et al., 2022) based on the GPT-3.5 model. Initially trained on a dataset of 570GB, ChatGPT supports input lengths of 4k and 16k, and demonstrates remarkable capabilities in human communication, boasting a substantial knowledge base, proficiency in reasoning through mathematical problems, accurate contextual understanding in multi-turn dialogues, alignment with human values to ensure safe use and following diverse user instructions. Subsequently, the introduction of GPT4 (OpenAI, [n. d.]) further extends the model’s capabilities, enabling the resolution of multi-modal tasks involving both image and text inputs and outputs. It is worthy noting that the maximum input window size is also doubled, which empowers the model to handle long context input. Benefiting from the cooperation with OpenAI, Microsoft release New Bing [^7] https://blogs.microsoft.com/wp-content/uploads/prod/sites/5/2023/02/The-new-Bing-Our-approach-to-Responsible-AI.pdf, which is a conversational agent based on GPT4, capabling to provide up-to-date information with the assistance of search engine.
Table 5. An overview of Dialogue Models based on Large Language Models listed in time order.
Dialogue Models Release Date Base Model # of Params Window size Trainig Tokens Language Preference Source Type BlenderBot3(Shuster and et al., 2022) Aug, 2022 OPT 3B/30B/175B 180B EN Open ChatGPT Dec, 2022 GPT3.5 (Brown et al., 2020) 175B 4k/16k EN Closed Bing Chat Feb, 2023 GPT4 (OpenAI, [n. d.]) 8k/32k EN Closed LLaMA-Chat (Touvron et al., 2023a) Feb, 2023 LLaMA 7B/13B/33B/65B 2k 1.4T EN Open Alpaca (Taori et al., 2023) Mar, 2023 LLaMA 7B 2k 1.4T EN Open Vicuna (Chiang and et al., 2023) Mar, 2023 LLaMA 7B/13B 2k 1.4T EN Open ChatGLM (Zeng et al., 2022) Mar, 2023 GLM 6B/12B 2k 1T CN Open Slack Mar, 2023 Claude 200k 2.0T EN Closed Bard Mar, 2023 PALM 540B EN Closed Falcon-Chat (Penedo and et al., 2023) Jun, 2023 Falcon 7.5B/40B/180B 2k 1.5T/1T/3.5T EN Open MPT-Chat (Lin and et al., 2023) Jun, 2023 MPT 7B/30B 8k 1T EN Closed InternLM-Chat (Team, 2023) Jun, 2023 InternLM 7B/20B 8k 2.3T CN Open Baichuan-Chat (Yang et al., 2023), Jun, 2023 Baichuan 7B/13B/53B 4k 1.2T CN Open (7B/13B) ChatGLM2 (Zeng et al., 2022) Jun, 2023 GLM 6B/12B 32k 1T CN Open LLaMA2-Chat (Touvron et al., 2023b) Jul, 2023 LLaMA2 7B/13B/34B/70B 4k 2.0T EN Open Atom-Chat Aug, 2023 LLaMA2 7B/13B 4k 1.0T CN Open Qwen-Chat (Bai and et al., 2023) Aug, 2023 QWen 7B/14B 8k 2.4T CN Open Baichuan2-Chat (Yang et al., 2023) Sep, 2023 Baichuan2 7B/13B 4k 2.6T CN Open ChatGLM3 Oct, 2023 GLM 6B 32k / CN Open
@@
On the other hand, unlike the closed-source nature of the GPT series models from OpenAI, Meta has embraced openness by releasing the LLaMA (Touvron et al., 2023a) and LLaMA2 (Touvron et al., 2023b), which are pre-trained on 1.4T data, offers four parameter configurations: 7B, 13B, 33B, and 65B. This provides a fresh opportunity for an open-source community, with corresponding dialogue models being consistently finetuned based on LLaMA. LLaMA-Chat, for example, finetunes the LLaMA model using instructional methods, aligning with the GPT-3 autoregressive paradigm. In addition, Alpaca (Taori et al., 2023) is finetuned by Stanford based on LLaMA-7B, using text-davinci-003 API and self-instruct techniques to create a prompt-response instruction dataset of 52K. Vicuna (Chiang and et al., 2023), based on LLaMA-13B, used supervised data from ShareGPT  [^8] https://sharegpt.com/, extending sequence length from 512 to 2048 during training. To further improve the window size and human values alignment with the public, LLaMA2 (Touvron et al., 2023b) employs grouped-query attention, extending input length from 2k to 4k and training data from 1.4T tokens to 2.0T tokens. Besides these two series, there are other series English LLMs such as Claude [^9] https://claude.ai/, Bard, Falcon, etc. These models typically adhere to an architecture resembling that of GPT-3, showcasing proficiency in handling multi-turn and complex dialogues, with different sizes of input window. Other commercial companies have also adopted similar training methods and subsequently introduced their own conversational large language models, such as MPT-Chat (Lin and et al., 2023) and Falcon-Chat (Penedo and et al., 2023).
@@
LLMs and DMs in Chinese. Besides English, there are also lots of impressive Chinese models. For example, ChatGLM (Du et al., 2022; Zhang and et al., 2023b), as the first open-source Chinese LLM, which is a bilingual model (Chinese and English) built on the general language model architecture – GLM 130B (Zhang and et al., 2023b). The open-sourced 6B version, ChatGLM2-6B, extends the context length from 2k to 32k tokens and employs multi-query attention for more efficient inference. Furthermore, they progress from ChatGLM and ChatGLM2 to ChatGLM3  [^10] https://github.com/THUDM/ChatGLM3/, which incorporates a broader and more diverse corpus of data during its fine-tuning process. Additionally, ChatGLM3 has demonstrated significant improvements in planning and reasoning, akin to language agents (Zeng et al., 2023; Sumers et al., 2023), thus reflecting substantial improvements in tool learning, code execution, database operation, knowledge graph search and reasoning. There are publicly available LLMs from other companies, such as Baichuan-Chat (Yang et al., 2023) and Baichuan2-Chat. The latter, Baichuan2-Chat, employs a method similar to Claude and is trained on over 2.6T tokens. These models cater to a wide spectrum of applications, each offering unique advantages stemming from their fine-tuning strategies and underlying architecture. Alibaba releases Qwen-7B-Chat aligns with human intent using a curated dataset, pre-trained on over 2.2T tokens with a context length of 2048. Sensetime publish InternLM-Chat, which is derived from supervised fine-tuning and reinforcement learning with human feedback on InternLM (Team, 2023), a multilingual model pre-trained on 2.3T tokens.
LLMs and DMs in Other Languages. Since most of existing LLMs naturally support multiple languages due to comprehensive corpus in different languages involved during pre-training, there is only few attempts to further finetune dialogue model for specific languages, especially for underrepresented languages. We summarize several such attempts for Korean, Sherkala, and more boarder languages at the bottom part of Table 6. Specifically, most studies directly finetune the large language model (i.e., QWen and LLaMA) on lots of multilingual corpus to get the multilingual versions. For examples, Suzume 8B (Devine, 2024) is a multilingual model fine-tuned on nearly 90,000 conversations using Llama 3, endowing it with Llama 3’s capabilities alongside enhanced multilingual support. In contrast, Panges-7B (Yue and et al., 2024) extends this functionality by incorporating multimodal capabilities in addition to multilingual features.
Table 6. Some of representative multilingual large language models.
Models Release Date Base Model # of Params Window size Trainig Tokens Language Preference Source Type Llama-2-ko-7B-Chat Dec 2023 LLaMA2 7B 2k / Korean Open Suzume 8B (Devine, 2024) May 2024 LLaMA3 8B 8k / Multilingual Open LLaMAX3-8B (Lu and et al., 2024) July 2024 LLaMA3 8B 8k / Multilingual Open Pangea-7B (Yue and et al., 2024) Oct 2024 QWen2 7B 32k / Multilingual Open Aya-Expanse (Dang and et al., 2024) Dec 2024 Command 8B 8k / Multilingual Open Babel-Chat (Zhao and et al., 2025) Feb 2025 QWen 9B/83B 131K / Multilingual Open Llama-3.1-Sherkala-8B-Chat (Koto and et al., 2025) Mar 2025 LLaMA3.1 8B 8k / Sherkala Open